The ratio between the present ages of Prakash and Vikas is 5 : 7 respectively. If the difference between Vikas’s present age and Prakash’s age after 6 years is 2, what is the total of their present ages ?
Answer by mananth(16946)

prakash to Vikas ratio = 5/7
multiply by common factor x
Prakash = 5x & Vikas = 7x
prakash after 6 years = 5x+6
7x-(5x+6)=2
7x-5x-6=2
2x=8
x=4
Prakash = 5*4 = 20
Vokas = 7*4 = 28
Sum = 20+28=48 years
m.ananth@hotmail.ca